What volume of NH3 gas in liters is produced at STP when 0.28 L of a 0.87 M NH4Cl solution reacts according to the above equation?


1
Expert's answer
2021-04-07T11:32:57-0400

NH3 + HCl -> NH4Cl

Molar Mass of NH4Cl = 53.491g/mol

0.8M NH4CL = 53.491 × 0.8 = 42.7928g

0.28L = 280g.= 280/42.7928 = 6.54mol

From the equation 1 mole of NH4Cl produces 1 mole of NH3

6.54 mol of 0.28L NH4Cl produces

1 × 22.4 L/mol × 6.54 mol

= 146L NH3